OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x - y = 13 4x + y = 13 4x + y = -13

OpenStudy (anonymous):

4*4 + (-3) + k = 0 16 - 3 + k = 0 13 + k = 0 k = -13. 4y + y - 13 = 0 Hope this helps!